OpenStudy (chaise):

If z1 and z2 are complex numbers, prove that: (z1 + (conjugate z1)(z2+ (conjugate z2) = z1z2 + (conjugate z1z2) + z1(conjugate z2) + (conjugate(z1(conjugate z2) I'm not sure if I should be writing this in a different format. I don't even think you'll be able to help me with this one.

OpenStudy (chaise):

\[(z _{1}+(conjugatez _{1})) (z _{2}+(conjugatez _{2}) = z _{1}z _{2}+(conjugate z _{1}z _{2}) + z _{1}(conjugatez _{2}+(conjugatez _{1}(conjugatez _{2}))\]

OpenStudy (anonymous):

Just foil it out.. \[\large (z_1 + \widetilde{z_1})(z_2 + \widetilde{z_2}) \]\[\large = z_1z_2 + \widetilde{z_1}z_2 + z_1\widetilde{z_2} + \widetilde{z_1}\widetilde{z_2}\]

OpenStudy (anonymous):

All goes to 0?

OpenStudy (chaise):

So once it has been foiled out - can we let: z1 = a+bi z1(conjugate)= a-bi z2 = c+di z2 (conjugate) = c-di? then do some algebra and manipulation of the two to make them equal?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Yep, that'll work.

OpenStudy (nilankshi):

good